OpenStudy (anonymous):

There are x feet in y miles. Which of the following is an expression for the number of feet in z miles? A. xz/y B. yz/x C. y/xz D. x/yz E. xyz

jimthompson5910 (jim_thompson5910):

hint: x feet = y miles x/y feet = 1 mile

OpenStudy (anonymous):

Can I get another hint...

OpenStudy (texaschic101):

I am sorry..I don't know this one....I will stand back and watch

jimthompson5910 (jim_thompson5910):

if 1 mile = x/y feet then z miles = ??? how do we go from 1 to z (what's one way)

jimthompson5910 (jim_thompson5910):

say 1 mile = 3 ft (obviously not true, but just go with it) then 2 miles = 6 ft (multiply both sides by 2) 3 miles = 9 ft (multiply both sides by 3) etc etc

OpenStudy (texaschic101):

and 4 miles = 16 ft and 5 miles =25 ft....I just don't know where the z comes in

jimthompson5910 (jim_thompson5910):

so if 1 mile = 3 ft, then multiplying both sides by z would give you z miles = 3z ft

jimthompson5910 (jim_thompson5910):

see how I did that?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

if 1 mile = x ft, then z miles = xz feet again I got this by multiplying both sides by z

OpenStudy (anonymous):

Would it be xz/y?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (texaschic101):

like I said..he is good

OpenStudy (anonymous):

Haha yeah :)

jimthompson5910 (jim_thompson5910):

it's not too bad really once you get used to the notation

OpenStudy (anonymous):

You explain it better than my teacher at my academy

jimthompson5910 (jim_thompson5910):

I'm glad I do (although I do with you had better teachers over there)

OpenStudy (anonymous):

Yeah its only my math teacher though.

jimthompson5910 (jim_thompson5910):

I gotcha

OpenStudy (anonymous):

THank you again :)

jimthompson5910 (jim_thompson5910):

sure thing